OpenStudy (anonymous):

what is the expression in standard form by expanding and collecting like terms 4(8m-7n)+6(3n-4m)

alones (alones):

Umm Hi there =] you have any answer choices?? ;3

OpenStudy (anonymous):

no they dont give me any

alones (alones):

Oh okay hun let me show oyu how to solve it k

OpenStudy (anonymous):

okay

alones (alones):

Start by distribute: =(4)(8m)+(4)(−7n)+(6)(3n)+(6)(−4m) =32m+−28n+18n+−24m Combine Like Terms: =32m+−28n+18n+−24m =(32m+−24m)+(−28n+18n) =8m+−10n <---- i got this one actually let me tag some one to check it ;3

alones (alones):

@zepdrix can you check this??

OpenStudy (anonymous):

yea thanx

alones (alones):

yw dear ;)

OpenStudy (anonymous):

okay